### v1

# Some remarks on the results derived by Ramy Takieldin and Patrick Solé (2025)

[…]
Some remarks on the results derived by
Ramy Takieldin and Patrick Solé (2025)
Varsha Chauhan [^1] Email address: vchauhan@maitreyi.du.ac.in  and Anuradha Sharma [^2] Email address: anuradha@iiitd.ac.in
Department of Mathematics, Maitreyi College, New Delhi 110021, India
Department of Mathematics, IIIT-Delhi, New Delhi 110020, India
In [4], the authors begin by presenting counterexamples to Theorems 5.5 and 5.6 of [3]. Additionally, they made an unsubstantiated claim that “these examples emphasize that imposing a condition on the dimension of the multi-twisted (MT) code or its dual to guarantee being LCD is impossible. For instance, the MT code given in Example 1 has dimension greater than $\min\{m_{i}\},$ while its dual has dimension smaller than $\min\{m_{i}\},$ and that in Example 2 has dimension equal to $\min\{m_{i}\}$, and all of which are not LCD ” (see page 4, lines 2 - 5 of the last paragraph).
The purpose of this note is to address a typographical error in the statements of Theorems 5.5 and 5.6 of [3], which can be corrected by replacing “$\Lambda$-multi-twisted code” with “$[\Lambda,\Lambda^{\prime}]$-multi-twisted code”. The proofs of Theorems 5.5 and 5.6 of [3] are correct. Moreover, Corollary 5.1 of [3] is correct and follows directly from Theorem 5.2 of [3] and Corollary 2.6 of Liu and Liu [2]. This correction also resolves the issue raised by the authors of [4] in Remark 3.1(b) of [1], where it is stated that the results derived in Section 5 of [3] can be similarly extended when the block lengths $m_{i}$'s are not necessarily coprime to $q,$ particularly the BCH type distance bound obtained in Theorem 5.3 of [3].
Another purpose of this note is to analyze and discuss the significance of the results derived in [4]. For this, we first state Theorem 7 of [4].
Theorem 7 of [4]. Let $\Lambda=\left(\lambda_{1},\lambda_{2},\ldots,\lambda_{\ell}\right)$, where $\lambda_{i}$ is a non-zero element of $\mathbb{F}_{q}$ for $1\leq i\leq\ell$. Let $\mathcal{C}$ (or $\mathcal{C}^{\perp}$) be a $\Lambda$-MT code of block lengths $\left(m_{1},m_{2},\ldots,m_{\ell}\right)$ generated by $\left\{\mathbf{g}_{1},\mathbf{g}_{2},\ldots,\mathbf{g}_{\rho}\right\}$. For $1\leq i\leq\ell$, let $g_{i}\left(x\right)=\mathrm{gcd}\left\{\left(x^{m_{i}}-\lambda_{i}\right),\pi_{i}\left(\mathbf{g}_{1}\right),\pi_{i}\left(\mathbf{g}_{2}\right),\ldots,\pi_{i}\left(\mathbf{g}_{\rho}\right)\right\}$. If
$\frac{x^{m_{1}}-\lambda_{1}}{g_{1}\left(x\right)},\frac{x^{m_{2}}-\lambda_{2}}{g_{2}\left(x\right)},\ldots,\frac{x^{m_{\ell}}-\lambda_{\ell}}{g_{\ell}\left(x\right)}$
are pairwise coprime polynomials in $\mathbb{F}_{q}[x]$, then $\mathcal{C}$ and $\mathcal{C}^{\perp}$ are direct sums of $\ell$ constacyclic codes, in particular
$\begin{split}\mathcal{C}&=\bigoplus_{i=1}^{\ell}\pi_{i}\left(\mathcal{C}\right)\\ \mathcal{C}^{\perp}&=\bigoplus_{i=1}^{\ell}\pi_{i}\left(\mathcal{C}\right)^{\perp}=\bigoplus_{i=1}^{\ell}\pi_{i}\left(\mathcal{C}^{\perp}\right).\end{split}$
Remark 1. The above theorem implies that when $\frac{x^{m_{1}}-\lambda_{1}}{g_{1}\left(x\right)},\frac{x^{m_{2}}-\lambda_{2}}{g_{2}\left(x\right)},\ldots,\frac{x^{m_{\ell}}-\lambda_{\ell}}{g_{\ell}\left(x\right)}$ are pairwise coprime polynomials in $\mathbb{F}_{q}[x]$, the $\Lambda$-multi-twisted code $\mathcal{C}$ (as defined in the above theorem) is a direct sum of the codes $\pi_{1}(\mathcal{C}),\pi_{2}(\mathcal{C}),\ldots,\pi_{\ell}(\mathcal{C}),$ where $\pi_{i}(\mathcal{C})$ is a $\lambda_{i}$-constacyclic code of length $m_{i}$ over $\mathbb{F}_{q}$ for $1\leq i\leq\ell.$ Further, the Euclidean dual code $\mathcal{C}^{\perp}$ is a direct sum of the dual codes $\pi_{1}(\mathcal{C})^{\perp},\pi_{2}(\mathcal{C})^{\perp},\ldots,\pi_{\ell}(\mathcal{C})^{\perp},$ where $\pi_{i}(\mathcal{C})^{\perp}$ is a $\lambda_{i}^{-1}$-constacyclic code of length $m_{i}$ over $\mathbb{F}_{q}$ for $1\leq i\leq\ell.$ From this, it follows that if $G_{i}$ is a generator matrix of the $\lambda_{i}$-constacyclic code $\pi_{i}(\mathcal{C})$ of length $m_{i}$ over $\mathbb{F}_{q},$ then the $\Lambda$-multi-twisted code $\mathcal{C}=\bigoplus\limits_{i=1}^{\ell}\pi_{i}\left(\mathcal{C}\right)$ of length $n=m_{1}+m_{2}+\cdots+m_{\ell}$ and block lengths $\left(m_{1},m_{2},\ldots,m_{\ell}\right)$ over $\mathbb{F}_{q}$ has a generator matrix
$G=\begin{bmatrix}G_{1}&0&0&\cdots&0\\ 0&G_{2}&0&\cdots&0\\ \cdots&\cdots&\cdots&\cdots&\cdots\\ \cdots&\cdots&\cdots&\cdots&\cdots\\ 0&0&0&\cdots&G_{\ell}\end{bmatrix}.$
This further implies that
$d(\mathcal{C})=\min\{d(\pi_{1}(\mathcal{C})),d(\pi_{2}(\mathcal{C})),\ldots,d(\pi_{\ell}(\mathcal{C}))\},$ (0.1)
where $d(\cdot)$ denotes the Hamming distance function. Thus, the Hamming distance (and consequently, the error-correction capability) of the $\Lambda$-multi-twisted code $\mathcal{C}$ cannot exceed the Hamming distances (and error-correction capabilities) of any of the constacyclic codes $\pi_{i}(\mathcal{C})$ for $1\leq i\leq\ell$. Furthermore, the code $\mathcal{C}$ has length $n=m_{1}+m_{2}+\cdots+m_{\ell}$, while the constacyclic code $\pi_{i}(\mathcal{C})$ has length $m_{i}$ for $1\leq i\leq\ell$. Consequently, the class of $\Lambda$-multi-twisted codes examined in Theorem 7 of [4] lacks significant relevance from a coding theory perspective. Notably, not all multi-twisted codes over finite fields and their dual codes are direct sums of constacyclic codes. Thus, the majority of results derived in [1] and Section 5 of [3] focus on multi-twisted codes and their special subclasses, which are not necessarily direct sums of constacyclic codes.
In Theorem 5.2 of [3], we obtained a generating set of the Euclidean dual code of a $\Lambda$-multi-twisted code of length $n=m_{1}+m_{2}+\ldots+m_{\ell}$ and block lengths $\left(m_{1},m_{2},\ldots,m_{\ell}\right)$ over $\mathbb{F}_{q}$ when the polynomials $x^{m_{1}}-\lambda_{1},x^{m_{2}}-\lambda_{2},\ldots,x^{m_{\ell}}-\lambda_{\ell}$ are pairwise coprime in $\mathbb{F}_{q}[x].$ Later, in Theorem 3.1 of [1] (resp. Theorem 3.2 of [1]), we showed that each $\Lambda$-multi-twisted code of length $n=m_{1}+m_{2}+\ldots+m_{\ell}$ and block lengths $\left(m_{1},m_{2},\ldots,m_{\ell}\right)$ over $\mathbb{F}_{q}$ has a normalized generating set (resp. unique nice normalized generating set), and further, in Theorem 3.3 of [1], we explicitly obtained a generating set of its dual code in terms of the normalized generating set of the $\Lambda$-multi-twisted code of length $n=m_{1}+m_{2}+\ldots+m_{\ell}$ and block lengths $\left(m_{1},m_{2},\ldots,m_{\ell}\right)$ over $\mathbb{F}_{q}$, where the polynomials $x^{m_{1}}-\lambda_{1},x^{m_{2}}-\lambda_{2},\ldots,x^{m_{\ell}}-\lambda_{\ell}$ are not necessarily pairwise coprime in $\mathbb{F}_{q}[x]$ and $m_{1},m_{2},\ldots,m_{\ell}$ are arbitrary positive integers, not necessarily coprime to $q.$ However, in [4], the authors only discussed Theorem 5.2 of [3] and did not discuss Theorem 3.3 of [1], thereby providing an incomplete and inaccurate representation of the existing results in this direction.
We next state Corollary 8 of [4].
Corollary 8 of [4]. Let $\Lambda=\left(\lambda_{1},\lambda_{2},\ldots,\lambda_{\ell}\right)$, where $0\neq\lambda_{i}\neq\lambda_{i}^{-1}$ for $1\leq i\leq\ell$. Let $\mathcal{C}$ (or $\mathcal{C}^{\perp}$) be a $\Lambda$-MT code of block lengths $\left(m_{1},m_{2},\ldots,m_{\ell}\right)$ generated by $\left\{\mathbf{g}_{1},\mathbf{g}_{2},\ldots,\mathbf{g}_{\rho}\right\}$. For $1\leq i\leq\ell$, let $g_{i}\left(x\right)=\mathrm{gcd}\left\{\left(x^{m_{i}}-\lambda_{i}\right),\pi_{i}\left(\mathbf{g}_{1}\right),\pi_{i}\left(\mathbf{g}_{2}\right),\ldots,\pi_{i}\left(\mathbf{g}_{\rho}\right)\right\}$. Assume
$\frac{x^{m_{1}}-\lambda_{1}}{g_{1}\left(x\right)},\frac{x^{m_{2}}-\lambda_{2}}{g_{2}\left(x\right)},\ldots,\frac{x^{m_{\ell}}-\lambda_{\ell}}{g_{\ell}\left(x\right)}$
are pairwise coprime polynomials in $\mathbb{F}_{q}[x]$, then $\mathcal{C}$ is LCD.
Remark 2. By Theorem 7 of [4], the code $\mathcal{C}$ (as defined in Theorem 7 of [4]) is LCD if and only if $\pi_{i}(\mathcal{C})$ is LCD for $1\leq i\leq\ell.$ By applying Corollary 2.6 of Liu and Liu [2], Corollary 8 of [4] follows immediately.
Theorem 9 of [4] is as stated below.
Theorem 9 of [4]. Let $\Lambda=\left(\lambda_{1},\lambda_{2},\ldots,\lambda_{\ell}\right)$, where $\lambda_{i}$ is a non-zero element of $\mathbb{F}_{q}$ for $1\leq i\leq\ell$. Let $\mathcal{C}$ (or $\mathcal{C}^{\perp}$) be a $\Lambda$-MT code of block lengths $\left(m_{1},m_{2},\ldots,m_{\ell}\right)$ generated by $\left\{\mathbf{g}_{1},\mathbf{g}_{2},\ldots,\mathbf{g}_{\rho}\right\}$. For $1\leq i\leq\ell$, let $g_{i}\left(x\right)=\mathrm{gcd}\left\{\left(x^{m_{i}}-\lambda_{i}\right),\pi_{i}\left(\mathbf{g}_{1}\right),\pi_{i}\left(\mathbf{g}_{2}\right),\ldots,\pi_{i}\left(\mathbf{g}_{\rho}\right)\right\}$. If
$\frac{x^{m_{1}}-\lambda_{1}}{g_{1}\left(x\right)},\frac{x^{m_{2}}-\lambda_{2}}{g_{2}\left(x\right)},\ldots,\frac{x^{m_{\ell}}-\lambda_{\ell}}{g_{\ell}\left(x\right)}$
are pairwise coprime polynomials in $\mathbb{F}_{q}[x]$, then $\mathcal{C}$ is LCD if and only if $g_{i}\left(x\right)$ is self-reciprocal coprime to $\frac{x^{m_{i}}-\lambda_{i}}{g_{i}\left(x\right)}$ for every $i$ with $\lambda_{i}^{2}=1$.
Remark 3. Note that Theorem 9 of [4] is a straightforward consequence of Theorem 7 of [4] and an easy extension of the main theorem derived in Yang and Massey [5] to $\lambda$-constacyclic codes over $\mathbb{F}_{q},$ where $\lambda=\lambda^{-1}$.
Remark 4. When $\frac{x^{m_{1}}-\lambda_{1}}{g_{1}\left(x\right)},\frac{x^{m_{2}}-\lambda_{2}}{g_{2}\left(x\right)},\ldots,\frac{x^{m_{\ell}}-\lambda_{\ell}}{g_{\ell}\left(x\right)}$ are pairwise coprime polynomials in $\mathbb{F}_{q}[x],$ Corollary 8 and Theorem 9 of [4] provided methods to construct LCD codes over finite fields. LCD codes over finite fields are primarily studied due to their application in constructing orthogonal direct-sum masking schemes for protecting sensitive data against fault-injection and side-channel attacks. The security threshold (against fault-injection and side-channel attacks) of the orthogonal direct-sum masking scheme based on an LCD code is equal to the Hamming distance of the LCD code. Now, when $\frac{x^{m_{1}}-\lambda_{1}}{g_{1}\left(x\right)},\frac{x^{m_{2}}-\lambda_{2}}{g_{2}\left(x\right)},\ldots,\frac{x^{m_{\ell}}-\lambda_{\ell}}{g_{\ell}\left(x\right)}$ are pairwise coprime polynomials in $\mathbb{F}_{q}[x],$ by Theorem 7 of [4], it follows that the code $\mathcal{C}$ (as defined in Theorem 7 of [4]) is LCD if and only if the $\lambda_{i}$-constacyclic code $\pi_{i}(\mathcal{C})$ is LCD for $1\leq i\leq\ell.$ Thus the security thresholds $S,$ $S_{1},S_{2},\ldots,S_{\ell}$ against fault-injection and side-channel attacks of the orthogonal direct-sum masking schemes constructed using the LCD codes $\mathcal{C},$ $\pi_{1}(\mathcal{C}),\pi_{2}(\mathcal{C}),\ldots,\pi_{\ell}(\mathcal{C})$ respectively, satisfy the following:
$S=d(\mathcal{C})\text{ and }S_{i}=d(\pi_{i}(\mathcal{C}))\text{~{} for ~{}}1\leq i\leq\ell,$
which, by (0.1), gives
$S=\min\{S_{1},S_{2},\cdots,S_{\ell}\}.$
That is, the security threshold of the orthogonal direct-sum masking scheme designed using the LCD code $\mathcal{C}$ can not exceed the security threshold of the orthogonal direct-sum masking schemes designed using any of the LCD constacyclic codes $\pi_{1}(\mathcal{C}),\pi_{2}(\mathcal{C}),\ldots,\pi_{\ell}(\mathcal{C}).$ Therefore, the LCD codes obtained in Corollary 8 and Theorem 9 of [4] hold limited practical significance.
In Corollary 3.1 of [1], we obtained the dimension of a $\rho$-generator multi-twisted code of block lengths $(m_{1},m_{2},\ldots,m_{\ell})$ over $\mathbb{F}_{q}$ by summing up the degrees of the diagonal polynomials in the normalized generating set of the code, where $\rho\geq 1$ is any integer and $m_{1},m_{2},\ldots,m_{\ell}$ are arbitrary positive integers, not necessarily coprime to $q$. In [4], the authors highlighted the computation of a normalized generating set for a multi-twisted code as a drawback of Corollary 3.1 of [1] used to determine the dimension of the multi-twisted code (see page 2, lines 27-29 of [4]), which is again an unsubstantiated claim, as the proof of Theorem 3.1 of [1] outlines an approach to find a normalized generating set of the multi-twisted code. Highlighting this drawback of Corollary 3.1 of [1], the authors, in Theorem 13 of [4], suggested an alternate approach to obtain the dimension of a $\rho$-generator multi-twisted code of block lengths $(m_{1},m_{2},\ldots,m_{\ell})$ over $\mathbb{F}_{q}$, where one needs to obtain the greatest common divisor of all $\ell\times\ell$ minors of its generator matrix, which is a $(\rho+\ell)\times\ell$ matrix over $\mathbb{F}_{q}[x].$ We believe that computing $\ell\times\ell$ minors of a generator matrix of a $\rho$-generator multi-twisted code of block lengths $(m_{1},m_{2},\ldots,m_{\ell})$ over $\mathbb{F}_{q}$ and then finding their greatest common divisor may be computationally hard for large values of $\ell$ and $\rho$ in general. In [4], the authors also claim that Theorem 13 of [4] generalizes Theorem 5.1(b) of [3] (see page 8, the first sentence in Section 4 and page 9, lines 4-5 of [4]), which is also not true, as Theorem 13 of [4] does not coincide with Theorem 5.1(b) of [3] when $\rho=1$.
In our opinion, several claims made by the authors in [4] are either factually incorrect or lack adequate substantiation, which may confuse the readers about the contributions of [1, 3]. Our remarks on the work [4] intend to provide the clarity and inform about the true contributions and findings of our research.